OpenStudy (anonymous):

Please help I will Medal! Write the expression shown as one scientific number. (8 x 10^2)(7.5 x 10^4) _______________________ 5 x 10^2 I honestly don't know how to do this. If someone could help and give me an explanation that would help a lot.

satellite73 (satellite73):

first multiply the two numbers \(8\times 7.5\) in the numerator

OpenStudy (anonymous):

ok

satellite73 (satellite73):

then compute \(10^2\times 10^4\) by adding the exponents

satellite73 (satellite73):

divide the result of \(8\times 7.5\) by \(5\)

OpenStudy (anonymous):

wait ok so 8 x 7.5 is 60 right?

satellite73 (satellite73):

yes now divide that by 5

OpenStudy (anonymous):

12

satellite73 (satellite73):

right so you have \[\frac{12\times 10^2\times 10^4}{10^2}\]

satellite73 (satellite73):

cancel the common factor of \(10^2\) top and bottom

OpenStudy (anonymous):

so I'll have 12 × 10^4 ?

satellite73 (satellite73):

yes one more step

OpenStudy (anonymous):

ok

satellite73 (satellite73):

since \(12>10\) this number is not written in scientific notation you have to put a decimal point between the \(1\) and the \(2\) in the number \(12\) and also add \(1\) to the exponent

OpenStudy (anonymous):

right ok so it would be 1.2 x 10^5 ?

OpenStudy (anonymous):

@satellite73

satellite73 (satellite73):

yes

OpenStudy (anonymous):

Thank you so much this really helped a lot!

satellite73 (satellite73):

yw

OpenStudy (anonymous):

@satellite73 Sorry to bother you again but could you explain really quick why I had to divide 60 by 5?